HIGH-RESOLUTION LASER SPECTROSCOPY OF THE A^{1}Σ^{+} − X^{1}Σ^{+} TRANSITION OF CaS

\underline{TODD C. MELVILLE} and JOHN A. COXON, Department of Chemistry, Dalhousie University, Halifax, NS, Canada B3H 4J3.

High-resolution laser excitation spectroscopy has been employed for the first time to examine the A^{1}Σ^{+} −X^{1}Σ^{+} system of CaS. Gaseous CaS was produced in a Broida-type oven by the reaction of calcium metal vapor with OCS. A Coherent 699-29 ring dye laser operating in single-frequency mode was used to investigate the rotational structure of the A−X transition. The present work extends the lower resolution work of Blues and Barrow ^{a} on the same system to higher vibrational levels and has allowed the indentification of several level crossings that occur in the A^{1}Σ^{+} state. The deperturbation analysis of the A − X transition will be discussed.

aTrans. Faraday Soc. 65, 646 (1969)